B. Less than 50? 1. Convert the raw score of 50 to a z-score. 2. Draw the normal curve and locate the given z-value or values at the base line of the curve. Then draw a vertical line through the given z-value or values and shade the required region.


1
Expert's answer
2021-05-14T11:46:25-0400

(a) We have that:

"\mu=42"

"\sigma=5"


z score is


"Z=\frac{x-\mu}{\sigma}=\frac{50-42}{5}=1.6"


(b) Normal curve is-